OpenStudy (e.cociuba):

***HELP****MEDAL****** x^2 + 1x +13 Factor it.

OpenStudy (dan815):

you can use completing the square method or the quadratic equation

OpenStudy (dan815):

google quadratic equation

OpenStudy (e.cociuba):

Umm thanks for ur help i guess.

OpenStudy (dan815):

x^2 + 1x +13 = completing square is like this (x^2 + 1x + 1/4) + 13 - 1/4 = 0 (x+1/2 )(x+1/2) +(13-1/4) = 0 (x+1/2) ^2 = 51/4 x = +/- sqrt(51/4) -1/2

OpenStudy (dan815):

that is completing square method u can use just quadratc equation if u dont wanna think about it

OpenStudy (e.cociuba):

Wow! Thanks. I got the answer already, I just wanted to check if it was right. Thanks!

OpenStudy (dan815):

for completing the square basically you look at that x^2 + 1x and say what number shud i put beside it to make it a perfect square but subtract that number of your already present constant number in this case you can put 1/4 beside it making it (x^2 + 1x + 1/4) +13 - 1/4 = (x+1/2)^2 +13-1/4

OpenStudy (dan815):

you can just looked up a video on how to complete the square on youtube it will explain nicely in 2 -3 min video

OpenStudy (e.cociuba):

I need help with this: (If u can help) Can you envision ways in which people in technology in underdeveloped countries, could benefit from inexpensive internet?

OpenStudy (dan815):

yes cheap education lol

OpenStudy (e.cociuba):

Haha hey tht's actually a good one. lol

OpenStudy (dan815):

the uses for internet are unlimited, it will set up frame work for business give the ability to a common citizen to join the export and import business with the whole world open up new job options on the internet anything u can dream of lmao

OpenStudy (e.cociuba):

I was thinking about communicating with relatives in other countries, better education etc.

OpenStudy (dan815):

think about it like this what is it that you cant do with internet?

OpenStudy (dan815):

but underdeveloped countries have a big problem with jobs and the internet gives them another platform of available jobs there all kinds of jobs on the internet from markettings to research and development

OpenStudy (e.cociuba):

I guess u can do almost about any thing u put ur mind to.

OpenStudy (e.cociuba):

Who ur a genius. Thanks!

OpenStudy (e.cociuba):

R u in college?

OpenStudy (dan815):

commerce and media are a big thing and being able to share ideas on how to progressively enhance the underdeveloped country with the new avaiable acess to different minds around the world that are well versed in engineering and government

OpenStudy (dan815):

yeah lol im in university xD

OpenStudy (dan815):

far from a genius lmao you will see what i mean once you get older :)

OpenStudy (e.cociuba):

Nice! And thanks again for ur help. I'm in college and High School at the same time. It's alot on my plate, but I guess u can do anything u put ur mind to.

OpenStudy (dan815):

oh that must be a lot of work, i wish you best of luck keep at it!

OpenStudy (e.cociuba):

Thanks and u too!

OpenStudy (dan815):

this is gonna sound kinda funny but to be honest if im ever stuck on problems or it gets too hard I just pray and God helps me :)

OpenStudy (e.cociuba):

Lol i'm a christian too! U?

OpenStudy (e.cociuba):

It's not funny at all! God Bless U!

OpenStudy (e.cociuba):

@dan815

OpenStudy (dan815):

yeah I am christian:)

OpenStudy (e.cociuba):

Good for u!!!